# When Do Transformers Learn Heuristics for Graph Connectivity?

### v1

@@ Abstract
Transformers often fail to learn generalizable algorithms, instead relying on brittle heuristics. Using graph connectivity as a testbed, we explain this phenomenon both theoretically and empirically. We consider a simplified Transformer architecture, the disentangled Transformer, and prove that an $L$-layer model has capacity to solve for graphs with diameters up to exactly $3^{L}$, implementing an algorithm equivalent to computing powers of the adjacency matrix. We analyze the training-dynamics, and show that the learned strategy hinges on whether most training instances are within this model capacity. Within-capacity graphs (diameter $\leq 3^{L}$) drive the learning of a correct algorithmic solution while beyond-capacity graphs drive the learning of a simple heuristic based on node degrees. Finally, we empirically demonstrate that restricting training data within a model’s capacity leads to both standard and disentangled transformers learning the exact algorithm rather than the degree-based heuristic.
@@ 1 Introduction
Large language models (LLMs) based on the Transformer architecture have demonstrated remarkable capabilities, yet their success is often shadowed by failures on tasks that demand robust, algorithmic reasoning. A growing body of evidence shows that, instead of learning generalizable algorithms, these models frequently rely on brittle shortcuts and spurious correlations that exploit statistical cues in the training data (Niven & Kao, 2019; Geirhos et al., 2020; Tang et al., 2023; Yuan et al., 2024; Zhou et al., 2024b; Ye et al., 2024). This shortcut reliance contributes to poor out-of-distribution (OOD) generalization, vulnerability to adversarial prompts, and unreliability on multi-step reasoning tasks (Zou et al., 2023; Deng et al., 2024; Li et al., 2024). Evidence spans domains: in natural language inference, models pick up lexical-overlap heuristics rather than syntactic reasoning (McCoy et al., 2019; Cosma et al., 2024); and in mathematical problem solving, strong in-distribution scores often fail to transfer as problem structure or size shifts (Saxton et al., 2019; Kao et al., 2024; Zhou et al., 2025). This motivates a foundational question:
@@
When and why do Transformers learn heuristics over verifiably correct algorithms,
even when the task admits an algorithmic solution?
To study when Transformers learn algorithms rather than shortcuts, we adopt graph connectivity as a controlled testbed. Connectivity offers a unique ground-truth solution: given an adjacency matrix $A$ with self-loops, reachability equals the transitive closure and is computable by classical dynamic programming (Warshall, 1962; Floyd, 1962), so the target is unambiguous. Connectivity is a fundamental algorithmic problem and is one of the most well-studied problems in terms of its complexity (Wigderson, 1992). It is a natural starting point since it sits low in the complexity hierarchy, undirected sstt connectivity is known to lie in deterministic logspace $\mathsf{L}$ (Reingold, 2008). Recent theory further shows that Transformers with depth growing logarithmically in input length can implement nontrivial parallel algorithms—including connectivity—via repeated squaring-style constructions (Merrill & Sabharwal, 2025). At the same time, connectivity admits simple heuristics based on global statistics such as node degrees or local density, which can be predictive on many graphs but are misleading on others. By combining an unambiguous algorithmic target, a complexity landscape tied to Transformer depth, and natural shortcut baselines, connectivity is a principled stress test of whether training yields multi-step algorithmic composition or superficial cues.
Our contributions. Despite theoretical expressivity guarantees, whether gradient descent can help Transformers find the algorithmic solution remains unknown. In this work, our preliminary experiments reveal a clear algorithm-heuristic tension on graph connectivity (see §3.3 and Figure 1): Transformer models achieve perfect in-distribution accuracy but fail to generalize. We consider a simplified Transformer architecture, the disentangled Transformer (Friedman et al., 2023; Nichani et al., 2024) to analyze when training recovers an algorithmic solution rather than a shortcut. Finally we empirically show that the same pattern transfers to standard Transformer models. We summarize our contributions below:
  1. 1.
@@
    Non-asymptotic capacity tied to diameter. Prior work establishes log-depth expressivity for connectivity: Transformer models require a depth of $L\in\Theta(\log(n))$ in the number of nodes $n$ (Merrill & Sabharwal, 2025). In Theorem 4.4, we prove a non-asymptotic capacity theorem that depends on instance difficulty, characterized by the graph diameter rather than only on the number of nodes $n$. Let $\mathrm{diam}(G)$ denote the maximum shortest-path distance between any two connected nodes. We show that an $L$-layer model solves connectivity on all graphs with $\operatorname{diam}(G)\leq 3^{L}$. We refer to $3^{L}$ as the model’s capacity. We complement this by proving a matching capacity bound of $3^{L}$, and we empirically validate the diameter-depth scaling by training both disentangled and standard Transformer models.
  2. 2.
    An algorithm-heuristic decomposition. We prove in Theorem 4.6 and §4.6 that if the model has certain symmetries such as being invariant to relabellings of the vertices of the graph, the learned weights for a disentangled Transformer are a superposition of an algorithmic and a heuristic channel. We empirically validate that trained models have this invariance property. The algorithmic channel is responsible for multi-hop composition or the computation of matrix powers of the adjacency matrix (via repeated squaring). The heuristic channel determines if two nodes are connected based on the degrees of the two nodes, and similar higher-order generalizations of the degree based on the local neighborhood of the two nodes.
  3. 3.
@@
    Training dynamics. Our analysis of the training dynamics reveals a sharp dichotomy driven by the data distribution. For graphs within the model’s capacity (diameter $\leq 3^{L}$), population gradients suppress the heuristic channel and favor the algorithmic channel that implements matrix powering (Theorem C.5). Conversely, when the distribution contains a significant share of beyond-capacity graphs (diameter $3^{L}$) the gradients instead strengthen the heuristic channel, promoting the simple degree-counting shortcut (Theorem C.9). This precise characterization hinges on our exact $3^{L}$ capacity bound; an asymptotic one, such as the $\mathcal{O}(\exp(L))$ result from Merrill & Sabharwal (2025), would not yield such clear predictive implications.
  4. 4.
@@
    The Data Lever. These theoretical insights point to a direct mitigation strategy we call the data lever: restricting the training data exclusively to within-capacity graphs. Our experiments in §5 confirm the effectiveness of this approach, showing that it boosts the algorithmic component and improves out-of-distribution robustness (Figure 4), and that these benefits transfer successfully to standard Transformer models (Figure 7).
With graph connectivity as a testbed, our results together pinpoint precise breaking points of Transformers and how the training data influences the learning of generalizable algorithmic components versus brittle heuristics. Our analysis yields a strategy to reduce dependence on heuristics, via the data lever, demonstrating that the theory also has some prescriptive power.
@@ 2 Related Work
Computational Complexity and Expressivity of Transformers.
Theoretical analyses aim to define what Transformers can and cannot compute. Although Transformers are universal approximators for continuous sequence-to-sequence functions (Yun et al., 2020), they also face sharp complexity-theoretic limits. Fixed-depth attention struggles with periodic or hierarchical patterns (Hahn, 2020), and standard Transformers are restricted to the complexity class $\mathsf{TC}^{0}$ (Merrill & Sabharwal, 2023), with hard-attention variants also confined to low-level circuit classes (Hao et al., 2022; Barceló et al., 2024). This computational power can be expanded. Allowing model depth to scale logarithmically with input length enables recognition of regular languages and solves graph connectivity (Merrill & Sabharwal, 2025), while chain-of-thought generation also strictly increases expressivity (Merrill & Sabharwal, 2024). Programmatic abstractions like $\mathsf{RASP}$ offer another lens, identifying which algorithms can be implemented in a length-generalizing way (Weiss et al., 2021; Zhou et al., 2023). Empirically for the graph connectivity problem, Fu et al. (2024b) shows frontier LLMs can reach almost perfect performance on small graphs and Saparov et al. (2025) shows transformer has greater difficulty in learning the task when graph size increases. Additionally, Sanford et al. (2024) proves that logarithmic depth is both necessary and sufficient for parallelizable graph tasks, with supporting GraphQA evidence.
Mechanistic Interpretability of Transformers.
A growing body of work reverse-engineers the algorithmic circuits that Transformers learn for tasks like copying, induction, and reasoning (Elhage et al., 2021; Olsson et al., 2022; Wang et al., 2022; Brinkmann et al., 2024). These can range from Fourier-style circuits for modular addition (Nanda et al., 2023; Zhou et al., 2024a) to Newton-like updates for in-context linear regression (Fu et al., 2024a). Researchers validate hypotheses by compiling programs into model weights (Lindner et al., 2023), decompiling models into code (Friedman et al., 2023), and using causal interventions to localize function (Chan et al., 2022; Meng et al., 2022; Yao et al., 2024; Chang et al., 2024). On the other hand, Wen et al. (2023) show that head-by-head mechanistic explanations can mislead: even on a simple parentheses task, transformers spread a stack-like computation across many parts of the network. Finally, theoretical work on inductive biases, like a preference for low-sensitivity functions, helps explain why models often favor robust heuristics over exact algorithms (Vasudeva et al., 2025).
@@ 3.1 Graph Connectivity Task
Let $G=(V,E)$ be a finite simple graph with $n$ vertices. We define the self-loop-augmented adjacency matrix $A\in\{0,1\}^{n\times n}$ as: $A_{i,j}=1$ if $\{v_{i},v_{j}\}\in E$ or $i=j$, and 0 otherwise.
This definition is equivalent to taking the standard adjacency matrix and adding the identity matrix ($A_{G}+I_{n}$). A key consequence is that the $(i,j)$-th entry of the matrix power $A^{k}$ counts the number of walks of length $k$ from $v_{i}$ to $v_{j}$. With self-loops, these walks may stay at the same vertex from one step to the next. Henceforth, “adjacency matrix” will refer to this self-loop-augmented version.
@@
Our learning objective is to learn models $\mathcal{M}:\{0,1\}^{n\times n}\rightarrow\mathbb{R}^{n\times n}$. For a graph $G$ with adjacency matrix $A$ and connectivity matrix $R$, if $\mathcal{M}$ satisfies $\mathcal{M}(A)]_{i,j}>0\Leftrightarrow R_{i,j}=1$, then we say $\mathcal{M}$ is perfect on graph $G$. We train on Erdős-Rényi $\mathsf{ER}(n,p)$ graphs: on $n$ vertices, each edge is present independently with probability $p$.
@@ 3.2 Transformer Architectures
$\mathsf{ReadIn}(X):=XW_{\mathrm{in}},\quad W_{\mathrm{in}}\in\mathbb{R}^{n\times d},\qquad\mathsf{ReadOut}(H):=HW_{\mathrm{out}}^{\top},\quad W_{\mathrm{out}}\in\mathbb{R}^{n\times d}.$
An $L$-layer single-head transformer model for graph connectivity acts as
$\mathsf{TF}^{L}_{\Theta}(A)\;:=\;\mathsf{ReadOut}\!\left(\mathsf{Transformers}^{L}\!\big(\mathsf{ReadIn}(\bar{A})\big)\right)\in\mathbb{R}^{n\times n}\;$
where $\mathsf{Transformers}^{L}$ is a standard pre-norm Transformer with self-attention and with no causal attention masks. There is no additional positional encoding since $I_{n}$ is already added to the input $A$ as the absolute positional encoding. A full definition can be found in Appendix A.
@@ 3.3 Preliminary Study
[…]
[width=]figures/prelim_study/graph.pdf
[…]
[width=]figures/prelim_study/model_accuracy.pdf
Figure 1: We train 2-layer Transformers models on Erdős-Rényi graphs. (Left) Visualization of input, target and model prediction of a sample graph. (Right) Although trained models are able to predict perfectly on every edge within distribution, they failed to generalize to out-of-distribution graphs such as graphs with two isolated chains or cliques.
We train 2-layer Transformer models on $\mathsf{ER}(n=20,p=0.08)$ graphs and test them on two out-of distribution datasets: (1) $\mathsf{2Chain}(n=20,k=10)$ graphs with $n$ nodes consisting two isolated chains each with $k$ nodes, and (2) $\mathsf{2Clique}(n=20,k=10)$ graphs with $n$ nodes consisting two isolated $k$-Cliques. We measure the performance of model $\mathcal{M}$ via a exact match accuracy on our graph distribution $\mathcal{G}$, i.e., the fraction of graphs on which $\mathcal{M}$ is perfect. Formally, it’s defined as $\mathsf{ExactMatchAcc}(\mathcal{M},\mathcal{G})=\operatorname{\mathbb{E}}_{G=(V,E)\in\mathcal{G}}\left[\prod_{v_{i},v_{j}\in V}\mathbf{1}\left\{[\mathcal{M}(A_{G})]_{i,j}=[R_{G}]_{i,j}\right\}\right].$
Transformers Fail to Generalize. As shown in Figure 1, the 2-layer Transformer model is able to achieve almost perfect exact match accuracy on the held-out set of the training distribution. However, it fails to learn an algorithmic solution which would transfer to other distributions. When the model is tested on the $\mathsf{2Chain}$ and $\mathsf{2Clique}$ distributions, its exact match accuracy falls to nearly zero, indicating over-fitted heuristics have dominated the model prediction. We repeat the experiments via extensive hyperparameter search and scaling up the number of layers, but all models fail to generalize. This motivates us to further investigate this generalization failure on why transformers prefer learning heuristics, when this happens during the training dynamics, and if models can be mitigated to learn actual algorithmic solutions instead of overfitting heuristics.
@@ 4.1 Disentangled Transformer
To understand the generalization failure in §3.3 theoretically, we pivot to a simplified disentangled Transformer which helps us with not only expressivity/capacity analysis in §4.2 but also with training dynamics analysis in §4.3. In the disentangled Transformer, each attention block appends its output as a new coordinate slice of the residual stream rather than summing, so the representation dimension grows with depth and the read/write pathways become traceable (Friedman et al., 2023; Nichani et al., 2024). This model is more than a theoretical convenience; it serves as a valid proxy for its standard counterpart. Nichani et al. (2024) show that any standard attention-only transformer can be re-expressed as a disentangled model by specializing attention to implement feature concatenation, and Chen et al. (2024) adopt this architecture precisely because it preserves the computations of interest while being markedly more amenable to theoretical analysis. We now formalize the model.
@@
$\textrm{{Input hidden state}}\qquad h_{0}$ $=[I_{n},A]\in\mathbb{R}^{n\times d_{0}}$ (1)
$\textrm{{Hidden states at layer $\ell$}}\qquad h_{\ell}$ $=\left[h_{\ell-1},\mathrm{Attn}\left(h_{\ell-1};W_{\ell}\right)\right]\in\mathbb{R}^{n\times d_{\ell}}$
$\textrm{{Output layer}}\qquad\mathsf{TF}^{L}_{\Theta}(A)$ $=h_{L}W_{O}^{\top}$
$\textrm{where~~}\quad\mathrm{Attn}\left(h_{\ell-1};W_{\ell}\right):=\frac{1}{n}\mathsf{ReLU}\left(h_{\ell-1}W_{\ell}h_{\ell-1}^{\top}\right)h_{\ell-1}$ (2)
We remark that $h_{\ell}\in\mathbb{R}^{n\times d_{\ell}}$ where $d_{\ell}=2^{\ell+1}n$ grows exponentially with respect to $\ell$.

### v2

# Transformers Provably Learn Algorithmic Solutions for Graph Connectivity, But Only with the Right Data

@@ Abstract
Transformers often fail to learn generalizable algorithms, instead relying on brittle heuristics. Using graph connectivity as a testbed, we explain this phenomenon both theoretically and empirically. We consider a simplified Transformer architecture, the Disentangled Transformer, and prove that an $L$-layer model can compute connectivity in graphs with diameters up to $3^{L}$, implementing an algorithm equivalent to computing powers of the adjacency matrix. By analyzing training dynamics, we prove that whether the model learns this strategy hinges on whether most training instances are within this model capacity. Within-capacity graphs (diameter $\leq 3^{L}$) drive the learning of the algorithmic solution while beyond-capacity graphs drive the learning of a simple heuristic based on node degrees. Finally, we empirically show that restricting training data to stay within a model’s capacity makes both standard and Disentangled Transformers learn the exact algorithm.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) based on the Transformer architecture have demonstrated remarkable capabilities, yet their success is often shadowed by failures on tasks that demand robust, algorithmic reasoning. A growing body of evidence shows that, instead of learning generalizable algorithms, these models frequently rely on brittle shortcuts and spurious correlations that exploit statistical cues in the training data (Niven and Kao, 2019; Geirhos et al., 2020; Tang et al., 2023; Yuan et al., 2024; Zhou et al., 2024d; Ye et al., 2025). This shortcut reliance contributes to poor out-of-distribution (OOD) generalization, brittleness under superficial input changes, and unreliability on multi-step reasoning tasks (Zou et al., 2023; Deng et al., 2024; Li et al., 2024; Mirzadeh et al., 2025). On deterministic tasks like shift ciphers, LLMs favor high-probability outputs over correct solutions (McCoy et al., 2023); in mathematical problem solving, strong in-distribution scores often fail to transfer as problem structure or size shifts (Saxton et al., 2019; Kao et al., 2024; Zhou et al., 2024b). This motivates a foundational question:
@@
When and why do Transformers learn heuristics over verifiably correct algorithms, even when the task admits an algorithmic solution?
To study this question, we adopt graph connectivity as a controlled testbed. Connectivity is a fundamental problem in computational complexity (Wigderson, 1992) that combines three properties suited to our analysis. First, it admits an unambiguous algorithmic solution: reachability equals the transitive closure and is computable via matrix powering (Warshall, 1962; Floyd, 1962). Second, recent theory shows that Transformers with depth $L=\Theta(\log n)$ can express this algorithm through matrix powering constructions (Merrill and Sabharwal, 2025), so the architecture is provably expressive enough. Third, connectivity admits simple heuristics based on node degrees that correlate with the correct answer on typical random graphs but fail on adversarial instances. This makes it possible to test whether training recovers the algorithm or settles for the shortcut.
Despite these expressivity guarantees, whether gradient descent actually finds the algorithmic solution remains open. Our experiments reveal that standard Transformers achieve perfect in-distribution accuracy on random graphs yet fail catastrophically on simple OOD instances such as two disjoint chains (see §3.3 and Figure 1). To understand both the failure mode and how to correct it, we analyze the Disentangled Transformer (Friedman et al., 2023; Nichani et al., 2024), a simplified architecture that is more amenable to theoretical analysis while preserving the essential computations. We summarize our contributions below.
An $L$-layer model solves connectivity up to diameter $3^{L}$, but no further.
@@
We prove tight bounds that characterize model capacity in terms of graph diameter rather than the number of nodes. Let $\operatorname{diam}(G)$ denote the maximum shortest-path distance between any two connected nodes. On the expressivity side, we show that an $L$-layer Disentangled Transformer can solve connectivity on all graphs with $\operatorname{diam}(G)\leq 3^{L}$ by implementing a matrix powering algorithm (Theorem 4.3). On the limitations side, we prove a matching upper bound: for any choice of nonnegative weights, there exists a graph with diameter $3^{L}+1$ on which the model fails (Theorem 4.5). Together, these results establish that $3^{L}$ is the maximum diameter an $L$-layer model can handle perfectly. We call this quantity the model’s capacity. We empirically validate this diameter-depth scaling on both disentangled and standard Transformers.
Learned weights decompose into algorithmic and heuristic channels.
We prove that under natural symmetry assumptions, specifically invariance to relabeling of graph vertices, the weights of a trained Disentangled Transformer decompose into two functionally distinct components (Theorem 4.7). The algorithmic channel preserves locality and implements multi-hop composition by computing powers of the adjacency matrix. The heuristic channel ignores graph structure and instead computes global statistics based on node degrees, predicting connectivity from whether two nodes both have high degree. We verify empirically that trained models satisfy the required symmetry, making this decomposition applicable to practical settings (§4.3).
Training dynamics select between channels based on the data distribution.
@@
Our analysis of the training dynamics reveals a sharp dichotomy driven by the data distribution. For graphs within the model’s capacity (diameter $\leq 3^{L}$), population gradients suppress the heuristic channel and favor the algorithmic channel that implements matrix powering (Theorem C.5). Conversely, when the distribution contains a significant share of beyond-capacity graphs (diameter $3^{L}$), the gradients instead strengthen the heuristic channel, promoting the simple degree-counting shortcut (Theorem C.9). This precise characterization hinges on our exact $3^{L}$ capacity bound; an asymptotic one, such as the $\mathcal{O}(\exp(L))$ result from Merrill and Sabharwal (2025), would not yield such clear predictive implications.
Transformers learns algorithmic solution with the right data.
@@
These theoretical insights point to a direct mitigation strategy we call the data lever: restricting the training data exclusively to within-capacity graphs. Our experiments in §5 confirm the effectiveness of this approach, showing that it boosts the algorithmic component and improves OOD robustness (Figure 4), and that these benefits transfer successfully to standard Transformer models (Figure 6).
Our results, validated on both Disentangled and standard Transformers (Figures 2, 9 and 6), provide a precise account of how Transformers learn algorithms given the right data.
Figure 1: We train 2-layer Transformer models on Erdős-Rényi graphs. (Top) Visualization of a graph, its input adjacency matrix, and the model’s (mis-)prediction of its connectivity. (Bottom) Although trained models are able to predict perfectly on every edge within distribution, they failed to generalize to out-of-distribution graphs such as graphs with two isolated chains or cliques.
@@ 2 Related Work
Theoretical analyses aim to define what Transformers can and cannot compute. Although Transformers are universal approximators for continuous sequence-to-sequence functions (Yun et al., 2020), they also face sharp complexity-theoretic limits. Fixed-depth attention struggles with periodic or hierarchical patterns (Hahn, 2020), and standard Transformers are restricted to the complexity class $\mathsf{TC}^{0}$ (Merrill and Sabharwal, 2023), with hard-attention variants also confined to low-level circuit classes (Hao et al., 2022; Barcelo et al., 2024). Allowing model depth to scale logarithmically with input length enables solving graph connectivity (Merrill and Sabharwal, 2025; Sanford et al., 2024). Chain-of-thought also increases expressivity (Merrill and Sabharwal, 2024; Feng et al., 2023) but we exclude it here to focus only on what the base architecture learns through gradient descent. Programmatic abstractions like $\mathsf{RASP}$ offer another lens, identifying which algorithms can be implemented in a length-generalizing way (Weiss et al., 2021; Zhou et al., 2024a). Empirically for the graph connectivity problem, Fu et al. (2024b) shows frontier LLMs can reach almost perfect performance on small graphs and Saparov et al. (2025) shows Transformers have greater difficulty in learning the task when graph size increases. Our matrix powering view is analogous to message passing in GNNs, where $L$ layers reach $L$-hop neighborhoods (Hamilton, 2020); attention achieves $3^{L}$ hops by tripling the range at each layer. We include more related work in the appendix §E.
@@ 3.1 Graph Connectivity Task
Let $G=(V,E)$ be a graph with $n$ vertices. We define the self-loop-augmented adjacency matrix $A\in\{0,1\}^{n\times n}$ as $A_{i,j}=1$ if $\{v_{i},v_{j}\}\in E$ or $i=j$, and 0 otherwise.
This definition is equivalent to taking the standard adjacency matrix and adding the identity matrix. A key consequence is that the $(i,j)$-th entry of the matrix power $A^{k}$ counts the number of walks of length $k$ from $v_{i}$ to $v_{j}$. With self-loops, these walks may stay at the same vertex from one step to the next. Henceforth, “adjacency matrix” will refer to this self-loop-augmented version.
@@
Our learning objective is to learn models $\mathcal{M}:\{0,1\}^{n\times n}\rightarrow\mathbb{R}^{n\times n}$. For a graph $G$ with adjacency matrix $A$ and connectivity matrix $R$, if $\mathcal{M}$ satisfies $\mathcal{M}(A)]_{i,j}>0\Leftrightarrow R_{i,j}=1$, then we say $\mathcal{M}$ is perfect on graph $G$. We train on Erdős-Rényi $\mathsf{ER}(n,p)$ graphs, i.e., graphs on $n$ vertices where each edge is present independently with probability $p$.
@@ 3.2 Transformer Architectures
$\mathsf{ReadIn}(X)$ $=XW_{\mathrm{in}},\quad W_{\mathrm{in}}\in\mathbb{R}^{n\times d},$
$\mathsf{ReadOut}(H)$ $=HW_{\mathrm{out}}^{\top},\quad W_{\mathrm{out}}\in\mathbb{R}^{n\times d}.$
An $L$-layer single-head transformer model acts as
$\mathsf{TF}^{L}_{\Theta}(A)\;:=\;\mathsf{ReadOut}\!\left(\mathsf{Transformers}^{L}\!\big(\mathsf{ReadIn}(A)\big)\right)\;$
where $\mathsf{Transformers}^{L}$ is a standard pre-norm Transformer with self-attention and with no causal attention masks. There is no additional positional encoding since $I_{n}$ is already added to the input $A$ as the absolute positional encoding. A full definition can be found in Definition A.1.
@@ 3.3 Preliminary Study
We train 2-layer Transformer models on $\mathsf{ER}(n=20,p=0.08)$ graphs and test them on two out-of-distribution datasets: (1) $\mathsf{2Chain}(n=20,k=10)$ graphs with $n$ nodes consisting of two isolated chains each with $k$ nodes, and (2) $\mathsf{2Clique}(n=20,k=10)$ graphs with $n$ nodes consisting of two isolated $k$-Cliques. We measure the performance of model $\mathcal{M}$ via an exact match accuracy on our graph distribution $\mathcal{G}$, i.e., the fraction of graphs on which $\mathcal{M}$ is perfect, or $\mathsf{ExactMatchAcc}(\mathcal{M},\mathcal{G})=\operatorname{\mathbb{E}}_{G=(V,E)\in\mathcal{G}}\left[\prod_{v_{i},v_{j}\in V}\mathbf{1}\left\{[\mathcal{M}(A_{G})]_{i,j}=[R_{G}]_{i,j}\right\}\right].$
Transformers Fail to Generalize. As shown in Figure 1, the 2-layer Transformer model is able to achieve almost perfect exact match accuracy on the held-out set of the training distribution. However, it fails to learn an algorithmic solution that transfers to other distributions. When the model is tested on the $\mathsf{2Chain}$ and $\mathsf{2Clique}$ distributions, its exact match accuracy falls to nearly zero, indicating over-fitted heuristics have dominated the model prediction. We repeat the experiments via extensive hyperparameter search and scaling up the number of layers, but all models fail to generalize. This motivates us to investigate why transformers prefer to learn brittle heuristics and how we can encourage them to learn algorithmic solutions instead.
@@ 4.1 Disentangled Transformer
To understand the generalization failure in §3.3 theoretically, we pivot to a simplified Disentangled Transformer; this helps us not only with expressivity/capacity analysis in §4.2 but also with training dynamics analysis in §4.3. In the Disentangled Transformer, each attention block appends its output as a new coordinate slice of the residual stream rather than summing, so the representation dimension grows with depth and the read/write pathways become traceable (Friedman et al., 2023; Nichani et al., 2024). This model serves as a reasonable proxy for its standard counterpart: Nichani et al. (2024) show that any standard attention-only Transformer can be re-expressed as a disentangled model by specializing attention to implement feature concatenation, and Chen et al. (2024) adopt this architecture precisely because it preserves the computations of interest while being markedly more amenable to theoretical analysis. We now formalize the model.
@@
$\textrm{{Input hidden state }}h_{0}=[I_{n},A]\in\mathbb{R}^{n\times d_{0}}$
-$\textrm{{$\ell$-th Hidden state }}h_{\ell}=\left[h_{\ell-1},\mathrm{Attn}\left(h_{\ell-1};W_{\ell}\right)\right]\in\mathbb{R}^{n\times d_{\ell}}$
$\textrm{{Output layer }}\mathsf{TF}^{L}_{\Theta}(A)=h_{L}W_{O}^{\top}$
where $\mathrm{Attn}\left(h_{\ell-1};W_{\ell}\right):=\frac{1}{n}\mathsf{ReLU}\left(h_{\ell-1}W_{\ell}h_{\ell-1}^{\top}\right)h_{\ell-1}$. We remark that $h_{\ell}\in\mathbb{R}^{n\times d_{\ell}}$ where $d_{\ell}=2^{\ell+1}n$ grows exponentially with respect to $\ell$.